# Investigating potential causes of Sepsis with Bayesian network structure learning

### v1

@@ 4.3 Knowledge-based structure and constraints
One of the objectives of this study is to compare the learnt graphs to a knowledge-based reference graph. Figure 2 presents the knowledge graph constructed based on literature review and clinical expertise as discussed in Section 2. Please note that, in the knowledge graph, ethnicity remains unconnected as no evidence was found in the literature relating ethnicity to Sepsis. However, we include this variable in the input data to investigate whether the graphs generated by the structure learning algorithms would be consistent with this prior knowledge.
Figure 2: The knowledge DAG constructed based on literature review and clinical expertise.
In addition to constructing the knowledge graph, we also produce a list of knowledge constraints consistent with the knowledge graph that could be incorporated into the learning process of the algorithms. Table 3 presents a set of directed edges which will serve as knowledge-based constraints. The decision on these constraints was established by asking our clinical expert (refer to Section 2) about their confidence in the relationships of these variables. Where the answer was 90% or higher, the constraint was generated.
@@
In addition to the directed constraints, we also explore possible temporal constraints, using the principle that later events cannot cause earlier events. These temporal constraints have been used in the past by, for example, [35]. Table 4 will define these restrictions for the learning algorithm through multiple prohibitions on directed edges. Temporal tiers are introduced to symbolise this. For example, no variables in Tiers greater than Tier 1 can cause variables in Tier 1. In addition, Tier 1 variables cannot be a cause of each other either. However, variables in other tiers can be causes of each other.
@@ 4.4 Evaluation
  • Investigating the relationship between the learnt graphs and the reference knowledge-based graph. This is achieved using the SHD score, which counts the number of differences, also referred to as the Hamming distance, between two graphs ([32]). In the software used in this study, which we discuss in Section 5, the SHD metric is computed by comparing differences between Completed Partially Directed Acyclic Graphs (CPDAGS). A CPDAG contains both directed and undirected edges that cannot be orientated given the observational data and represents a set of DAGs that belong to the same Markov equivalence class.
  • Counting the number of independent graphical fragments, or disjoint subgraphs, produced by each algorithm. This is important because information flow is not possible between independent graphical fragments, which is undesirable since we consider the input data to consist of related variables (with the exception of ethnicity).
@@ 5 Results
Structure learning was done using the algorithms specified in Section 4, implemented in the R package bnlearn ([36]; [37]). All of the algorithms used have hyperparameters that impact the output they create. As with previous relevant studies, we have employed the algorithms using their default hyperparameter settings, as it remains unclear in the literature how to best tune the hyperparameters in different settings or how to systematically compare results across different types of algorithms with different kinds of hyperparameters.
@@ 5.1 Structure learning performance
Table 5 summarizes the evaluation metrics for each of the structure learning algorithms, as well as the average structure obtained through the model averaging process and the knowledge-based structure.
The results show that the constraint-based methods and MMHC did not produce a single graphical fragment, unlike the score-based algorithms HC and Tabu, the hybrid H2PC, and the model averaging graph which produced a single graphical fragment that connects all risk factors, which is partly explained by the high number of edges these graphs contain. It is important to remember that in the knowledge-based graph ethnicity is not connected to the rest of the variables as no evidence was found in the literature to consider it a risk factor of Sepsis.
@@
In terms of the number of free parameters and edges, we can see that the knowledge-based graph only generates a higher number of free parameters than the structures generated by MMHC and PC-stable. This is not a surprise given that both of these graphs have few edges; generally, less dense graphs generated fewer free parameters. However, this is only a general principle, as we can see how complex the structure generated by Inter-IAMB is with a small number of edges. To be more specific, there is an association between the number of free parameters and the number of edges. This is expected since more edges tend to lead to a greater number of parents per node, which can have a dramatic effect on the number of parameters. Differences across the algorithms might be explained by a small number of variables that contain many categories and the way each algorithm incorporates these variables in the structures they generate.
The knowledge-based graph offers the worst result with respect to the BIC score. This means that the knowledge-based graph might not be complex enough to capture the interrelationships of variables in the data, and some important dependencies might have been missed. This is not surprising since the algorithms are designed to maximise BIC (or some other objective score), and we would generally not expect a knowledge graph to have a better score than the highest-scoring graph an algorithm discovers across the search space of graphs visited by the algorithm.
@@
Overall, there seem to be two groups of algorithms with respect to the number of edges discovered. The first group consists of the two score-based learning algorithms, the hybrid H2PC, and the model averaging structure. This group has a high number of edges. On the other hand, the two constraint-based algorithms and hybrid MMHC discovered a relatively low number of edges. Most of the variables used in this study refer to the presence or absence of medical diagnosis or procedures, these are binary variables. Interestingly, this type of data structure paired with large samples, such as the one in this study, seems to generate a large number of edges in score-based methods.
BIC scores, which represent a score that balances model fitting with model dimensionality, follow a similar pattern. Constraint-based algorithms are not designed to maximise BIC scores, which means it is natural to expect that score-based algorithms (or hybrid algorithms that employ score-based solutions) produce better BIC scores, as evidenced in the results. Although there are some differences when considering the LL instead of the BIC score, these differences are not substantial.
@@
Interestingly, only Diabetes $\rightarrow$ Number of Diagnoses appears in the knowledge-based graph. While this highlights considerable disagreements between knowledge and structures learnt from data, this does not necessarily imply that the different graphs are not reasonably accurate. For example, the 1st and 2nd edges from the list above present an interesting finding in that these edges are deemed clinically sensible. Still, our clinical expert would not have placed a higher than 90% confidence that these were causal relationships. This highlights the need to consider both expertise and data-driven approaches to derive new understanding.
Furthermore, there are 8 edges, representing around 3% of all the edges learnt by the algorithms, that are learnt by 5 out of the 6 algorithms. Two of these edges appear in the knowledge-based graph, i.e., Age $\rightarrow$ Alcohol dependence and Age $\rightarrow$ Number of Diagnosis. The algorithms learnt Alcohol dependence $\rightarrow$ Sex instead of Sex $\rightarrow$ Alcohol dependence.
Figure 3 presents a heatmap that highlights the edges that appear two or more times in the graphs learnt by the six different structure learning algorithms. We observe that the following variables had the highest numbers of direct effects (child nodes):
@@
The distribution for direct causes (parent nodes) seems less skewed than the distribution for direct effects (child nodes), shown in Figure 4, and there were no variables with large numbers of direct causes.
@@ 5.2 Structure learning with knowledge-based constraints
Table 6 summarises the evaluation metrics for each structure learning algorithm and the average structure when they incorporate the knowledge-based constraints detailed in Tables 3 and 4. Table 6 also includes the information for the knowledge-based structure.
The results show that the constraints caused the SHD scores to decrease, as expected since the constraints are guiding algorithms towards the knowledge-based graph. Moreover, the number of edges increased for all algorithms. On the other hand, the number of free parameters presents an unclear pattern since it increased for some algorithms and decreased for others. The BIC score decreased for all the structures except for the one generated by PC Stable. This might be due to increasing the complexity of a less dense structure and because constraint-based methods do not optimise BIC scores.
Table 6: Summary of Results
@@
With the knowledge constraints enforced on all six algorithms, we obtain the model-averaging graph in Figure 5. We can see that the parents of Sepsis are:
@@
Figure 5: The structure obtained through model averaging across the six algorithms, with the knowledge constraints imposed on the structure learning process.
@@ 5.3 Predictive validation of Sepsis
In the field of BNs, validation methods used in structure learning usually focus on the fit of the data to the whole graph; e.g., the Log-Likelihood score. While the global probability distribution is important, we are also naturally interested in the ability of the model to predict Sepsis specifically. From the perspective of predicting Sepsis, this structure was evaluated using the common confusion matrix, accuracy, sensitivity, and specificity indicators. The predictive validation strategy involved 10-fold cross-validation, where in every rotation, 90% of data was used to parametrise (train) the CBN and 10% was used for testing. The metrics presented below are an average calculated using the test data for each of the 10-fold. These metrics were used in the field of BNs by [6] to evaluate their BN model of Sepsis and [39] to evaluate their BN model of hypertension.
The R package caret ([40]) was used to calculate the confusion matrix, accuracy, sensitivity, and specificity. The results of these indicators are calculated by judging that an episode will include a Sepsis diagnosis if the probability of Sepsis is higher than 3.56% (average of Sepsis cases in train data). This was deemed a natural threshold that could be easily adapted for other datasets. The accuracy in making predictions for our model was 69.6%, sensitivity was 76.5%, and specificity was 69.4%. These results are generally in line with the results of the models from the literature review discussed in Section 1. The confusion matrix is shown in Table 7 below:
@@
However, some of the models reviewed had the more ambitious objective of early detection of Sepsis ([5], [8], [9]). The input data partly explains the fact that our model has similar performance to these models. The model presented in this paper relies on data available for commissioning purposes only. In other words, the BN models learnt in this study are trained on a subset of the data used by comparable studies. Given the data limitations and the performance metrics obtained, the predictions of the presence of Sepsis this model can make are reasonably accurate.
Figure 6 shows a smooth and convex ROC curve arching towards the upper left corner of the plot. This is the plot we would like to see in a good predictive model since the top left corner of the plot would be the ideal model. The Area Under the Curve (AUC) is 0.8 which indicates that there is an 80% chance that the model will be able to distinguish between Sepsis and no Sepsis. This is a further indication of the good performance of our model.
@@ 5.4 Causal Inference and Intervention
Simulating interventions in CBNs refers to manipulating the state of a variable in a model. In this study, this is equivalent to, for example, a policy intervention to reduce alcohol consumption. Unlike assessments on predictive accuracy, estimating the effect of intervention does not require that we perform the train-test data split, implying that the results discussed in this subsection are based on the entire training dataset.
@@
These results estimate the potential benefits of preventing these factors for the occurrence of Sepsis. Of course, there might be other benefits of this prevention not considered here. But these results show the relevance of modelling the interrelations of these risk factors in understanding how information is transmitted across the system to enable the discovery of possible interventions to reduce Sepsis cases.
6 Conclusion

### v2

@@ 4.3 Knowledge-based structure and constraints
One of the objectives of this study is to compare the learnt graphs to a knowledge-based reference graph. Figure 2 presents the knowledge graph constructed based on literature review and clinical expertise, as discussed in Section 2. Please note that, in the knowledge graph, ethnicity remains unconnected as no evidence was found in the literature linkingethnicity to Sepsis. Nonetheless, we include this variable in the input data to investigate whether the graphs generated by the structure learning algorithms would be consistent with this prior knowledge. In Figure 2, this structure is divided into subgraphs to facilitate analysis. The addition of aggregator nodes facilitates visualisation by preventing edges from becoming tangled.
Figure 2: The knowledge DAG constructed based on literature review and clinical expertise separated with subgraphs.
In addition to constructing the knowledge graph, we also produce a list of knowledge constraints consistent with it that could be incorporated into the algorithms’ learning process. Table 3 presents a set of directed edges that will serve as knowledge-based constraints. We decided on these constraints by asking our clinical expert (refer to Section 2) about their confidence in the relationships of these variables. Where the answer was 90% or higher, the constraint was generated.
@@
In addition to the directed constraints, we explore possible temporal constraints, using the principle that later events cannot cause earlier events. These temporal constraints have been used in the past by, for example, [35]. Table 4 will define these restrictions for the learning algorithm through multiple prohibitions on directed edges. Temporal tiers are introduced to symbolise this. For example, no variables in Tiers greater than Tier 1 can cause variables in Tier 1. In addition, Tier 1 variables cannot be a cause of each other either. This means edges were forbidden within Tier 1; that is, between variables in Tier 1. However, variables in other tiers can be caused by each other; that is, edges between variables within other tiers were allowed.
@@ 4.4 Evaluation
  • Investigating the relationship between the learnt graphs and the reference knowledge-based graph. This is achieved using the SHD score, which counts the number of differences, also called the Hamming distance, between two graphs ([32]). In the software used in this study, which we discuss in Section 5, the SHD metric is computed by comparing differences between Completed Partially Directed Acyclic Graphs (CPDAGs). A CPDAG contains both directed and undirected edges that cannot be orientated given the observational data and represent a set of DAGs that belong to the same Markov equivalence class.
  • Counting the number of independent graphical fragments, or disjoint subgraphs, produced by each algorithm. This is important because information flow is not possible between independent graphical fragments, which is undesirable since we consider the input data to consist of related variables (except ethnicity).
@@ 5 Results
Structure learning was done using the algorithms specified in Section 4, implemented in the R package bnlearn ([36]; [37]). All of the algorithms used have hyperparameters that impact the output they create. As with previous relevant studies, we have employed the algorithms using their default hyperparameter settings, as it remains unclear in the literature how to best tune the hyperparameters in different settings or how to systematically compare results across different types of algorithms with different kinds of hyperparameters. Table 5 below shows the default bnlearn parameters:
Table 5: Default parameters for selected structure-learning algorithms in bnlearn. The actual test (e.g., mi vs. cor) is chosen based on data type (discrete vs. continuous), where mi is mutual information, restart is the number of random restarts, setting strict to TRUE leads to a sparser network by enforcing weaker associations, and max hyperparameters set to inf indicate no constraints in the number of conditioning sets or the number of iterations in neighbouring graphs visits in constraint-based and score-based learning respectively.
Algorithm & Function Default Score/Test Key Default Parameters
Score-Based
Tabu (tabu()) BIC tabu = 10, max.iter = Inf, optimized = TRUE
Hill-Climbing (HC) (hc()) BIC restart = 0, max.iter = Inf, optimized = TRUE
Constraint-Based
PC-stable (pc.stable()) mi (discrete) or cor (continuous) alpha = 0.05, strict = TRUE, max.sx = Inf
Inter-IAMB (inter.iamb()) mi (discrete) or cor (continuous) alpha = 0.05, strict = TRUE, max.sx = Inf
Hybrid (Constraint + Score)
MMHC (mmhc()) Restrict: mi, $\alpha=0.05$; Maximize: BIC restrict = "mmpc", maximize = "hc", restart = 0, max.iter = Inf, optimized = TRUE
H2PC (h2pc()) mi (discrete) or cor (continuous) alpha = 0.05, strict = TRUE, max.sx = Inf, restart = 0, max.iter = Inf, optimized = TRUE
@@ 5.1 Structure learning performance
Table 6 summarises the evaluation metrics for each structure learning algorithm, the average structure obtained through the model averaging process, and the knowledge-based structure.
The results show that the constraint-based methods and MMHC did not produce a single graphical fragment, unlike the score-based algorithms HC and Tabu, the hybrid H2PC, and the model averaging graph, which produced a single graphical fragment that connects all risk factors, which is partly explained by the high number of edges these graphs contain. It is important to remember that in the knowledge-based graph, ethnicity is not connected to the rest of the variables as no evidence was found in the literature to consider it a risk factor of Sepsis.
@@
Regarding the number of free parameters and edges, we can see that the knowledge-based graph only generates a higher number of free parameters than the structures generated by MMHC and PC-stable. This is not a surprise given that both graphs have few edges; generally, less dense graphs generate fewer free parameters. However, this is only a general principle, as we can see how complex the structure generated by Inter-IAMB is with a small number of edges. To be more specific, there is an association between the number of free parameters and the number of edges. This is expected since more edges tend to lead to a greater number of parents per node, which can dramatically affect the number of parameters. Differences across the algorithms might be explained by a small number of variables containing many categories and how each algorithm incorporates these variables in the structures they generate.
The knowledge-based graph yields the worst result with respect to the BIC score. This means that the knowledge-based graph might not be complex enough to capture the interrelationships of variables in the data, and some important dependencies might have been missed. The fact that the knowledge graph produces a lower objective score, however, is not surprising. This is because the algorithms are designed to maximise BIC (or some other objective score), and we would generally not expect a manually constructed knowledge graph to have a better score than the highest-scoring graph an algorithm discovers.
@@
Overall, there seem to be two groups of algorithms with respect to the number of edges discovered. The first group consists of the two score-based learning algorithms, the hybrid H2PC, and the model averaging structure. This group has a high number of edges. On the other hand, the two constraint-based algorithms and hybrid MMHC discovered a relatively low number of edges. Most of the variables used in this study refer to the presence or absence of medical diagnosis or procedures; these are binary variables. Interestingly, this type of data structure paired with large samples, such as the one in this study, seems to generate a large number of edges in score-based methods.
BIC scores, representing a score that balances model fitting with model dimensionality, follow a similar pattern. Constraint-based algorithms are not designed to maximise BIC scores, so it is natural to expect score-based algorithms (or hybrid algorithms that employ score-based solutions) to produce better BIC scores, as evidenced in the results. Although there are some differences when considering the LL instead of the BIC score, these differences are not substantial.
@@
Interestingly, only Diabetes $\rightarrow$ Number of Diagnoses appears in the knowledge-based graph. While this highlights considerable disagreements between knowledge and structures learnt from data, this does not necessarily imply that the different graphs are not reasonably accurate. For example, the 1st and 2nd edges from the list above present an interesting finding: these edges are deemed clinically sensible. Still, our clinical expert would not have placed a higher than 90% confidence that these were causal relationships. This highlights the need to consider both expertise and data-driven approaches to derive new understanding.
Furthermore, there are eight edges, representing around 3% of all the edges learnt by the algorithms, that are learnt by 5 out of the six algorithms. Two of these edges appear in the knowledge-based graph, i.e., Age $\rightarrow$ Alcohol dependence and Age $\rightarrow$ Number of Diagnosis. The algorithms learnt Alcohol dependence $\rightarrow$ Sex instead of Sex $\rightarrow$ Alcohol dependence.
Figure 3 presents a heatmap highlighting the edges that appear two or more times in the graphs learnt by the six different structure learning algorithms. We observe that the following variables had the highest numbers of direct effects (child nodes):
@@
The distribution for direct causes (parent nodes), shown in Figure 4, seems less skewed than the distribution for direct effects (child nodes), and no variables had large numbers of direct causes.
@@ 5.2 Structure learning with knowledge-based constraints
Table 7 summarises the evaluation metrics for each structure learning algorithm and the average structure when they incorporate the knowledge-based constraints detailed in Tables 3 and 4. Table 7 also includes information on the knowledge-based structure.
The results show that the constraints caused the SHD scores to decrease, as expected, since the constraints are guiding algorithms towards the knowledge-based graph. Moreover, the number of edges increased for all algorithms. On the other hand, the number of free parameters presents an unclear pattern since it increased for some algorithms and decreased for others. The BIC score decreased for all the structures except for the one generated by PC Stable. This might be due to increasing the complexity of a less dense structure and because constraint-based methods do not optimise BIC scores.
Table 7: Summary of results when algorithms incorporate knowledge-based constraints
@@
With the knowledge constraints enforced on all six algorithms, we obtain the model-averaging graph in Figure 5. Figure 5 shows this structure split using aggregator nodes to prevent edges from becoming tangled, facilitating analysis. We can see that the parents of Sepsis are:
@@
Figure 5: The structure obtained through model averaging across the six algorithms, with the knowledge constraints imposed on the structure learning process. The figure is using aggregator nodes to facilitate analysis.
@@ 5.3 Predictive validation of Sepsis
In the field of BNs, validation methods used in structure learning usually focus on the fit of the data to the whole graph; e.g., the Log-Likelihood score. While the global probability distribution is important, we are also naturally interested in the ability of the model to predict Sepsis specifically. From the perspective of predicting Sepsis, this structure was evaluated using the common confusion matrix, accuracy, sensitivity, and specificity indicators. The predictive validation strategy involved 10-fold cross-validation, where in every rotation, 90% of data were used to parametrise (train) the CBN and 10% were used for testing. The metrics presented below are an average calculated using the test data for each of the 10-fold. These metrics were used in the field of BNs by [39] to evaluate their BN model of hypertension.
The R package caret ([40]) was used to calculate the confusion matrix, accuracy, sensitivity, and specificity. The results of these indicators are calculated by judging that an episode will include a Sepsis diagnosis if the probability of Sepsis is higher than 3.56% (average of Sepsis cases in train data). This was deemed a natural threshold that could be easily adapted to other datasets. The accuracy in making predictions for our model was 69.6%, sensitivity was 76.5%, and specificity was 69.4%. These results generally align with the results of the models from the literature review discussed in Section 1. The confusion matrix is shown in Table 8 below:
@@
However, some of the models reviewed had a more ambitious objective: early detection of Sepsis ([5], [8], [9]). The input data partly explains why our model performs similarly to these models. The model presented in this paper relies on data available for commissioning purposes only. In other words, the BN models learnt in this study are trained on a subset of the data used by comparable studies. Given the data limitations and the performance metrics obtained, the predictions of the presence of Sepsis this model can make are reasonably accurate.
Figure 6 shows a smooth and convex ROC curve arching towards the upper left corner of the plot. This is the plot we would like to see in a good predictive model since the top left corner of the plot would be the ideal model. The Area Under the Curve (AUC) is 0.8, indicating an 80% chance that the model will be able to distinguish between Sepsis and no Sepsis. This is a further indication of good model performance.
@@
In addition to our proposed approach, several studies have developed models to predict Sepsis or enable its early detection. These studies employed a wide range of techniques, including traditional regression models and BNs. However, all of these studies had access to broader clinical information than our dataset, which was restricted to billing data and did not include detailed clinical data such as biomarkers.
Table 9 presents the predictive performance of these models from the literature alongside our knowledge-based and averaged-structure models. Where available, we report commonly used performance metrics for classification tasks, including the Area Under the Curve (AUC), accuracy, sensitivity, and specificity. Since the studies did not explicitly name their models, we have referenced them by the first author of each work. Notably, [8] reports multiple metrics because their model was evaluated on several datasets.
It is important to note that these comparisons are not based on identical datasets. The models differ significantly in data type, sample size, and clinical settings. Therefore, while these metrics provide insight into the relative performance of our models and those in the literature, they should be interpreted cautiously regarding direct comparisons.
The results indicate that the knowledge-based model, while designed to capture known causal relationships, performs poorly at predictive accuracy compared to the model-averaging BN model described in this study and other data-driven models in the literature. This is consistent with the finding that the knowledge graph produces the lowest objective BIC score. However, this outcome is not entirely surprising, as knowledge graphs constructed by domain experts do not maximise data fitting by design but rather aim to capture causal relationships that models maximising predictive accuracy may fail to do so. As a result, it is often the case that knowledge-based graphical structures will miss data patterns that algorithms may recover. On the other hand, the average structure, which also incorporates domain knowledge, performs comparably to other models in the literature despite being trained on a weaker dataset. Given these findings, the best approach for future work may be to combine the strengths of both methods by combining the causal insights of knowledge graphs with data-driven techniques.
Table 9: Comparison of predictive performance (AUC, accuracy, sensitivity, and specificity) across relevant models from the literature and those presented in this study. The models are listed by the first author’s name as no specific model names were provided in the source papers. Entries left blank indicate metrics not reported in that study. ”KB Model” refers to the knowledge-based model, and ”Avg-Structure” refers to the averaged BN model.
Reference Model / Author AUC Accuracy Sensitivity Specificity
[5] Yee et al. (2019) 0.81
[6] Gupta et al. (2020) 0.84
[7] Haug & Ferraro (2016) 0.86 0.76
[8] Pawar et al. (2019) 0.75, 0.76, 0.78 0.80, 0.89, 0.82
[9] Yang et al. (2020) 0.97 0.99 0.75
KB Model (This study) 0.58 0.82 0.31 0.84
Avg-Structure (This study) 0.80 0.70 0.77 0.69
[…]
@@ 5.4 Causal Inference and Intervention
Simulating interventions in CBNs involves manipulating the state of a variable in a model. In this study, this is equivalent to, for example, a policy intervention to reduce alcohol consumption. Unlike assessments of predictive accuracy, estimating the effect of interventions does not require performing the train-test data split, implying that the results discussed in this subsection are based on the entire training dataset.
@@
These results estimate the potential benefits of preventing these factors for the occurrence of Sepsis. Of course, there might be other benefits of this prevention not considered here. However, these results show the relevance of modelling the interrelations of these risk factors in understanding how information is transmitted across the system to enable the discovery of possible interventions to reduce Sepsis cases.
5.4.1 Actionable Policy Recommendations
In light of these findings, we highlight the following evidence-based policy recommendations that have the potential to reduce COPD, Alcohol dependence, and Diabetes, thereby reducing the incidence of Sepsis:
  • Investing in smoking cessation programmes. For example, providing subsidised access to nicotine replacement therapies.
  • Screening for alcohol misuse and offering referral pathways to specialised addiction services. Increasing educational programmes to address alcohol misuse before it develops into dependence.
  • Expanding lifestyle modification programmes addressing balanced nutrition, regular physical activity, and weight management for patients with diabetes.
  • Consider increasing taxation on tobacco and alcohol products.
These recommendations not only aim to reduce COPD, Alcohol dependence, and Diabetes, but also contribute to lowering the burden of Sepsis and other related comorbidities. Investing in targeted interventions enables healthcare systems to better realise broader benefits for population health. Structural models such as those considered in this study enable improved modelling of the interrelations of risk factors and could identify and better estimate the impact of these interventions and their resulting broader benefits.
6 Conclusions